# Non-Abelian interference of topological edge states

### v1

@@ III.2 Non-Abelian topological transport
where $\theta_{d}=\int_{0}^{T}\left[E_{+,\text{II}}(t)-E_{-,\text{II}}(t)\right]dt$ is the accumulated dynamical phase difference. The second equality leverages the relationships presented in Eq. (48), which are associated with the non-Abelian braiding process $\tau_{1}^{-1}\tau_{2}^{-1}$. Notably, this result shows that particles injected initially at the left end of Chain $B$ can be transferred to the two ends of Chain $A$ with a continuously tunable ratio (0 to 1), governed by the relative dynamical phase $\theta_{d}$. Similarly, for the initial state $|L\rangle_{B}$, we observe the symmetric behavior. Furthermore, when the initial state is placed at one end of the other chains, the results exhibit cyclic mutual transport, both of which are summarized as follows
@@
The neglected global phase differs for each pair of edge states, yet the relative dynamical phases $\theta_{d}$ for each pair during one period are the same: $\int_{0}^{T}\left[E_{+,\text{I}}(t)-E_{-,\text{I}}(t)\right]dt=\int_{0}^{T}\left[E_{+,\text{II}}(t)-E_{-,\text{II}}(t)\right]dt=\int_{0}^{T}\left[E_{+,\text{III}}(t)-E_{-,\text{III}}(t)\right]dt$. When considering non-Abelian transport governed by the braiding operator $\tau_{2}^{-1}\tau_{1}^{-1}$, the cyclic mutual transport takes a distinct form as follows:
$|1_{A/B/C}\rangle$ $\rightarrow\cos\frac{\theta_{d}}{2}|1_{C/A/B}\rangle-i\sin\frac{\theta_{d}}{2}|L\rangle_{B/C/A},$ (70)
@@
Figure 6: Non-Abelian topological transport. Density distribution $\langle\hat{n}_{j}\rangle$ as a function of the evolution time $t$ for the braiding operator $\tau_{1}^{-1}\tau_{2}^{-1}$ (left) and $\tau_{2}^{-1}\tau_{1}^{-1}$ (right). The initial state is $|1_{B}\rangle$, with parameters set as $L=14$, $w=1$, $v_{0}=0.5$, $\eta=0.2$, and $T=2663$ (corresponding to $\theta_{d}=3\pi$).
We present the density distribution $\langle\hat{n}_{j}\rangle$ as a function of the evolution time $t$. Here, we select the initial state as $|1_{B}\rangle$, with the system parameters set to period $T=2663$ and chain length $L=14$ (corresponding to $\theta_{d}=3\pi$). Specifically, as shown in Fig. 6, the final density distribution localizes at the right end of Chain $C$ for the braiding operator $\tau_{1}^{-1}\tau_{2}^{-1}$, while it localizes at the right end of Chain $A$ for $\tau_{2}^{-1}\tau_{1}^{-1}$. It intuitively demonstrates the non-Abelian nature of topological transport, characterized by an operation-sequence-dependent outcome.
Actually, there has been extensive and comprehensive research on band braiding and non-Abelian state transfer in non-Hermitian systems [57, 58, 59, 60]. However, in these works, the demonstration of non-Abelian properties is often reliant on quasi-static processes, which imposes limitations on their practical applications. In this work, we overcome this limitation by implementing non-Abelian braiding protected by a time-dependent non-Hermitian symmetry, rather than by the Hamiltonian itself. Looking forward, we can build on these works to extend our results to systems with more braiding branches, explore their relevance to practical quantum computation, and develop a more comprehensive theoretical framework.
@@ III.3 Non-Abelian Hong-Ou-Mandel interference
Consider two identical bosons injected into the two ends of Chain $B$, with their initial two-boson state given by $|1_{B},L_{B}\rangle$. When subjected to an adiabatic evolution period governed by the braiding operator $\tau_{1}^{-1}\tau_{2}^{-1}$, these bosons transform according to Eq. 67. Specifically, the initial state $|1_{B},L_{B}\rangle$ evolves into $i\sin\theta_{d}(|1_{C},1_{C}\rangle+|L_{C},L_{C}\rangle)/\sqrt{2}+\cos\theta_{d}|1_{C},L_{C}\rangle$. We choose $\theta_{d}=\frac{2n+1}{2}\pi$ ($n\in\mathbb{Z}$), which corresponds to a 50:50 ratio; under this condition, the antibunching term vanishes. In this case, the state simplifies to the two-particle NOON state
@@
neglecting a global phase factor. As a maximally entangled state between the two ends of Chain $C$, this spatial NOON state fundamentally arises from the well-known HOM interference. Notably, the HOM interference exhibits non-Abelian nature; when subjected to adiabatic evolution governed by the braiding operator $\tau_{2}^{-1}\tau_{1}^{-1}$, the bosons transform according to Eq. 70, resulting in the final NOON state belonging to Chain AA
@@
Figure 7: Non-Abelian HOM interference. (a) Density distribution $\langle\hat{n}_{j}\rangle$ as a function of evolution time $t$ for the braiding operators $\tau_{1}^{-1}\tau_{2}^{-1}$ (left) and $\tau_{2}^{-1}\tau_{1}^{-1}$ (right). (b) Two-particle correlation $\Gamma_{q,r}$ at initial and final times for the same braiding operators $\tau_{1}^{-1}\tau_{2}^{-1}$ (left) and $\tau_{2}^{-1}\tau_{1}^{-1}$ (right)]. The initial state is $|1_{B},L_{B}\rangle$, with parameters set as $L=14$, $w=1$, $v_{0}=0.5$, $\eta=0.2$, and $T=2220$ (corresponding to $\theta_{d}=2.5\pi$).
To intuitively illustrate the non-Abelian nature of HOM interference (i.e., operation-sequence-dependent outcomes), we present the density distribution $\langle\hat{n}_{j}\rangle$ over evolution time $t$ during adiabatic evolution in Fig. 7(a). We initialize the system with the product state $|1_{B},L_{B}\rangle$, fixing system parameters to period $T=2220$ and chain length $L=14$ (corresponding to $\theta_{d}=2.5\pi$). As shown in Fig. 7(a), this leads to a clear contrast: the final density distribution is uniformly distributed at Chain $C$’s two ends for $\tau_{1}^{-1}\tau_{2}^{-1}$ (left), while it localizes at Chain $A$’s two ends for $\tau_{2}^{-1}\tau_{1}^{-1}$ (right). This directly reflects the non-Abelian nature.
To characterize the correlation properties of the entanglement in the generated two-particle NOON state, we compute the two-particle correlation $\Gamma_{q,r}=\langle\hat{b}_{q}^{{\dagger}}\hat{b}_{r}^{{\dagger}}\hat{b}_{r}\hat{b}_{q}\rangle$ and plot it in Fig. 7(b). At the initial time $t=0$, the correlation function shows an uncorrelated antibunching pattern localized on Chain $B$. Following one adiabatic period, a distinct transition occurs: for the braiding operator $\tau_{1}^{-1}\tau_{2}^{-1}$, the pattern evolves into a correlated bunching distribution at Chain $C$’s two ends (left); for $\tau_{2}^{-1}\tau_{1}^{-1}$, it localizes at Chain $A$’s two ends (right). Notably, the fidelity between the final state and the target state ($|\Psi_{{\rm NOON},C}\rangle$ or $|\Psi_{{\rm NOON},A}\rangle$) reaches 0.995 in both cases. Thus, we successfully realize the spatially entangled NOON state via non-Abelian HOM interference.
@@ IV SUMMARY AND DISCUSSION
In this article, we have systematically investigated the braiding of topological edge state pairs protected by dual symmetries, namely time-independent inversion symmetry and time-dependent interchain symmetry, based on coupled SSH chains. Specifically, we observed the twisted energy spectrum and braiding structure of these edge state pairs, which originate from the swapping property of the eigenvalues of the time-dependent interchain symmetry operator. Furthermore, we achieved symmetry-protected adiabatic evolution and realized high-fidelity, ratio-tunable topological transport from one end of a chain to the two ends of another chain. Through non-Abelian topological transport and HOM interference of edge states, we demonstrate the key signatures of non-Abelian braiding: specifically, the swapping of edge state pairs and the order dependence of braiding outcomes. Moreover, we point out that the symmetry-protected non-Abelian braiding can be extended to bulk states.

### v2

@@ III.2 Non-Abelian topological transport
where $\theta_{d}=\int_{0}^{T}\left[E_{+,\text{II}}(t)-E_{-,\text{II}}(t)\right]dt$ is the accumulated dynamical phase difference. The second equality leverages the relationships presented in Eq. (48), which are associated with the non-Abelian permutation process $\mathcal{M}_{1}\mathcal{M}_{2}$. Notably, this result shows that particles injected initially at the left end of Chain $B$ can be transferred to the two ends of Chain $A$ with a continuously tunable ratio (0 to 1), governed by the relative dynamical phase $\theta_{d}$. Similarly, for the initial state $|L_{B}\rangle$, we observe the symmetric behavior. Furthermore, when the initial state is placed at one end of the other chains, the results exhibit cyclic mutual transport, both of which are summarized as follows
@@
The neglected global phase differs for each pair of edge states, yet the relative dynamical phases $\theta_{d}$ for each pair during one period are the same: $\int_{0}^{T}\left[E_{+,\text{I}}(t)-E_{-,\text{I}}(t)\right]dt=\int_{0}^{T}\left[E_{+,\text{II}}(t)-E_{-,\text{II}}(t)\right]dt=\int_{0}^{T}\left[E_{+,\text{III}}(t)-E_{-,\text{III}}(t)\right]dt$. When considering non-Abelian transport governed by the permutation operator $\mathcal{M}_{2}\mathcal{M}_{1}$, the cyclic mutual transport takes a distinct form as follows:
$|1_{A/B/C}\rangle$ $\rightarrow\cos\frac{\theta_{d}}{2}|1_{C/A/B}\rangle-i\sin\frac{\theta_{d}}{2}|L_{C/A/B}\rangle,$ (70)
@@
Figure 6: Non-Abelian topological transport. Density distribution $\langle\hat{n}_{j}\rangle$ as a function of the evolution time $t$ for the permutation operator $\mathcal{M}_{1}\mathcal{M}_{2}$ (left) and $\mathcal{M}_{2}\mathcal{M}_{1}$ (right). The initial state is $|1_{B}\rangle$, with parameters set as $L=14$, $w=1$, $v_{0}=0.5$, $\eta=0.2$, and $T=2663$ (corresponding to $\theta_{d}=3\pi$).
We present the density distribution $\langle\hat{n}_{j}\rangle$ as a function of the evolution time $t$. Here, we select the initial state as $|1_{B}\rangle$, with the system parameters set to period $T=2663$ and chain length $L=14$ (corresponding to $\theta_{d}=3\pi$). Specifically, as shown in Fig. 6, the final density distribution localizes at the right end of Chain $C$ for the permutation operator $\mathcal{M}_{1}\mathcal{M}_{2}$, while it localizes at the right end of Chain $A$ for $\mathcal{M}_{2}\mathcal{M}_{1}$. It intuitively demonstrates the non-Abelian nature of topological transport, characterized by an operation-sequence-dependent outcome.
Actually, there has been extensive and comprehensive research on state permutations, band braiding, and non-Abelian state transfer in non-Hermitian systems [67, 68, 69, 70, 71, 72]. However, in these works, the demonstration of non-Abelian properties is often reliant on quasi-static processes, which imposes limitations on their practical applications. In this work, we overcome this limitation by implementing non-Abelian permutation protected by a time-dependent non-Hermitian symmetry, rather than by the Hamiltonian itself. Looking forward, we can build on these works to extend our results to systems with more state permutations or braiding branches, explore their relevance to practical quantum computation, and develop a more comprehensive theoretical framework.
@@ III.3 Non-Abelian Hong-Ou-Mandel interference
Consider two identical bosons injected into the two ends of Chain $B$, with their initial two-boson state given by $|1_{B},L_{B}\rangle$. When subjected to an adiabatic evolution period governed by the permutation operator $\mathcal{M}_{1}\mathcal{M}_{2}$, these bosons transform according to Eq. 67. Specifically, the initial state $|1_{B},L_{B}\rangle$ evolves into $i\sin\theta_{d}(|1_{C},1_{C}\rangle+|L_{C},L_{C}\rangle)/\sqrt{2}+\cos\theta_{d}|1_{C},L_{C}\rangle$. We choose $\theta_{d}=\frac{2n+1}{2}\pi$ ($n\in\mathbb{Z}$), which corresponds to a 50:50 ratio; under this condition, the antibunching term vanishes. In this case, the state simplifies to the two-particle NOON state
@@
neglecting a global phase factor. As a maximally entangled state between the two ends of Chain $C$, this spatial NOON state fundamentally arises from the well-known HOM interference. Notably, the HOM interference exhibits non-Abelian nature; when subjected to adiabatic evolution governed by the permutation operator $\mathcal{M}_{2}\mathcal{M}_{1}$, the bosons transform according to Eq. 70, resulting in the final NOON state belonging to Chain AA
@@
Figure 7: Non-Abelian HOM interference. (a) Density distribution $\langle\hat{n}_{j}\rangle$ as a function of evolution time $t$ for the permutation operators $\mathcal{M}_{1}\mathcal{M}_{2}$ (left) and $\mathcal{M}_{2}\mathcal{M}_{1}$ (right). (b) Two-particle correlation $\Gamma_{q,r}$ at initial and final times for the same permutation operators $\mathcal{M}_{1}\mathcal{M}_{2}$ (left) and $\mathcal{M}_{2}\mathcal{M}_{1}$ (right)]. The initial state is $|1_{B},L_{B}\rangle$, with parameters set as $L=14$, $w=1$, $v_{0}=0.5$, $\eta=0.2$, and $T=2220$ (corresponding to $\theta_{d}=2.5\pi$).
To intuitively illustrate the non-Abelian nature of HOM interference (i.e., operation-sequence-dependent outcomes), we present the density distribution $\langle\hat{n}_{j}\rangle$ over evolution time $t$ during adiabatic evolution in Fig. 7(a). We initialize the system with the product state $|1_{B},L_{B}\rangle$, fixing system parameters to period $T=2220$ and chain length $L=14$ (corresponding to $\theta_{d}=2.5\pi$). As shown in Fig. 7(a), this leads to a clear contrast: the final density distribution is uniformly distributed at Chain $C$’s two ends for $\mathcal{M}_{1}\mathcal{M}_{2}$ (left), while it localizes at Chain $A$’s two ends for $\mathcal{M}_{2}\mathcal{M}_{1}$ (right). This directly reflects the non-Abelian nature.
To characterize the correlation properties of the entanglement in the generated two-particle NOON state, we compute the two-particle correlation $\Gamma_{q,r}=\langle\hat{b}_{q}^{{\dagger}}\hat{b}_{r}^{{\dagger}}\hat{b}_{r}\hat{b}_{q}\rangle$ and plot it in Fig. 7(b). At the initial time $t=0$, the correlation function shows an uncorrelated antibunching pattern localized on Chain $B$. Following one adiabatic period, a distinct transition occurs: for the permutation operator $\mathcal{M}_{1}\mathcal{M}_{2}$, the pattern evolves into a correlated bunching distribution at Chain $C$’s two ends (left); for $\mathcal{M}_{2}\mathcal{M}_{1}$, it localizes at Chain $A$’s two ends (right). Notably, the fidelity between the final state and the target state ($|\Psi_{{\rm NOON},C}\rangle$ or $|\Psi_{{\rm NOON},A}\rangle$) reaches 0.995 in both cases. Thus, we successfully realize the spatially entangled NOON state via non-Abelian HOM interference.
@@ IV SUMMARY AND DISCUSSION
In this article, we have systematically investigated the permutation of topological edge state pairs protected by dual symmetries, namely time-independent inversion symmetry and time-dependent interchain symmetry, based on coupled SSH chains. Specifically, we observed the twisted energy spectrum and braiding structure of these edge state pairs, which originate from the swapping property of the eigenvalues of the time-dependent interchain symmetry operator. Furthermore, we achieved symmetry-protected adiabatic evolution and realized high-fidelity, ratio-tunable topological transport from one end of a chain to the two ends of another chain. Through non-Abelian topological transport and HOM interference of edge states, we demonstrate the key signatures of non-Abelian permutation: specifically, the swapping of edge state pairs and the order dependence of permutation outcomes. This symmetry-protected non-Abelian transport and interference can be directly generalized to scenarios with more than three pairs of edge states by designing appropriate symmetry operators and Hamiltonians [65]. Additionally, considering internal degrees of freedom such as spin may eliminate the need for next-nearest coupling required for the permutation of multiple pairs of edge states. Moreover, we point out that the symmetry-protected non-Abelian permutation can be extended to bulk states.
@@ Acknowledgements.
Data availability
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.